OpenStudy (anonymous):

An observer in a small boat is floating on the surface of a large lake with clear, but contaminated water with index of refraction n = 1.38. Nearby is a wooden post emerging from the water which is embedded in the lake bottom, a distance h = 7.20 meters below the water's surface. What is the maximum distance from the post at which the observer would be able to see the bottom of the post?

OpenStudy (shamim):

the answer will be 6.847m

OpenStudy (shamim):

i wanna explain my procedure to solve this problem

OpenStudy (shamim):

\[n=\frac{ 1 }{ \sin \theta _{c} }\]

OpenStudy (shamim):

here n=refractive index of water

OpenStudy (shamim):

\[\theta _{c}=critical \angle\]

OpenStudy (shamim):

\[or, 1.38=\frac{ 1 }{ \sin \theta _{c} }\]

OpenStudy (shamim):

\[or, \sin \theta _{c}=\frac{ 1 }{ 1.38 }\]

OpenStudy (shamim):

\[\theta _{c}=\sin^{-1} \frac{ 1 }{ 1.38 }\]

OpenStudy (shamim):

\[or, \theta _{c}=0.8105084862\]

OpenStudy (shamim):

now

OpenStudy (shamim):

\[\tan \theta=\frac{ 7.20 }{ d }\]

OpenStudy (shamim):

here 7.20m is the length of the woden post d= distance frm the post to observer

OpenStudy (shamim):

\[or, d=\frac{ 7.20 }{ \tan \theta _{c} }\]

OpenStudy (shamim):

\[or, d=\frac{ 7.20 }{ \sin.8105084862 }=6.847m\]

OpenStudy (shamim):

if u need more explanation then u can ask me